1. 4c-3=63 ?
2.3p+85=2p-10
3.6h+7=7h-5
4.14k-1=
5.7e-4=2e+13
1
Expert's answer
2012-09-14T08:25:53-0400
1. 4c-3=63
4с=63+3=66
с=66/4=33/2

2.3p+85=2p-10
3p-2p=-85-10
p=-95

3.6h+7=7h-5
7-5=7h-6h
h=2